OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

Y = mx + B m being the slope B is the y-intercept

OpenStudy (anonymous):

y = mx +b m =-6 x=-2 y=-3 solve for b -3=-6*(-2) +b

OpenStudy (anonymous):

y + 3 = -6(x + 2)

OpenStudy (anonymous):

Easy its y + 3 = -6(x + 2

OpenStudy (anonymous):

@kmeds16 @theredhead1617 don't just post the answer, explain to the questioner :}

OpenStudy (anonymous):

Sometimes they don't want you to. some times they just want the answer. fast.

OpenStudy (anonymous):

yeah i know, but it's in the rules :/

OpenStudy (anonymous):

don't want you getting banned and you not knowing why.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

didn't you read the terms and conditions? ;]